write an equation in slope-intercept form of the line that is parallel to the graph of the equation and passes through the given point
y=3x+6; (4,7)

Answer by mananth(16949)

y=3x+6; (4,7)
slope m = 3
the parallel line will also have a slope of 3
...
find b from y = mx +b
plug m & ( 4,7)
7=3*4+b
7=12+b
b=-5
...
The equation is
y = 3x-7